# Entanglement and apparent thermality in simulated black holes

### v1

@@ 1 Surface gravity of Gullstrand-Painlevé metric
Consider the (1+1)-dimensional Gullstrand-Painlevé metric line element given in Eq. (10), which is the line element of the (1+1)-dimensional Schwarzschild metric expressed in Gullstrand-Painlevé coordinates $x^{\mu}=(t,x)$. Unlike the Schwarzschild metric, the Gullstrand-Painlevé metric does not possess a coordinate singularity at $v^{2}=u^{2}$ and, therefore, is valid as both an internal and external metric. Here, the couplings $u$ and $v$ are independent of $t$, but not necessarily of $x$; therefore, the metric is stationary, though not static [21]. As the Gullstrand-Painlevé metric is stationary, it admits a Killing vector $\xi^{\mu}=\delta^{\mu}_{t}=(1,0)$ [57], which is time-like in the region $u^{2}>v^{2}$ and satisfies the Killing equation
$\xi_{\mu;\nu}+\xi_{\nu;\mu}=0.$ (C.1)
The vector $\xi_{\mu}=g_{\mu\nu}\xi^{\nu}=(1-v^{2}/u^{2},-v/u^{2})$ is also a Killing vector that satisfies the Killing equation. These Killing vectors can be used to find the surface gravity $\kappa$ of the black hole space-time, as defined by a distant observer, by evaluating
$\xi^{\mu}\xi_{\nu;\mu}=\xi^{\mu}\left(\xi_{\nu,\mu}-\xi_{\gamma}\tilde{\Gamma}^{\gamma}_{\nu\mu}\right)=\kappa\xi_{\nu},$ (C.2)
at the event horizon $x_{h}$ of the black hole [4, 61].
Expanding the Einstein summations over $\mu$ and $\gamma$, and making use of the fact that $\xi_{\nu,t}=\forall\nu$, $\xi^{t}=1$, and $\xi^{x}=0$, Eq. (C.2) for the surface gravity can be expressed as
$\xi_{x}\tilde{\Gamma}^{x}_{\nu t}-\xi_{t}\tilde{\Gamma}^{t}_{\nu t}=\kappa\xi_{\nu},$ (C.3)
@@
$\frac{1}{2}\xi_{t}g^{tt}g_{tt,x}-\frac{1}{2}\xi_{x}g^{xt}g_{tt,x}=\kappa\xi_{x}.$ (C.4)
Substitution of the relevant values for the components of the covariant and contravariant Gullstrand-Painlevé metric tensors, which are given in Eq. (A.19), into the above equation and evaluating the resultant expression at the event horizon gives the following surface gravity:
$\kappa=\pm\frac{1}{u^{2}}\frac{du}{dx}\Big.\Big|_{x=x_{h}}-\frac{1}{u^{2}}\frac{dv}{dx}\Big.\Big|_{x=x_{h}},$ (C.5)
as defined for an observer that is located an infinite distance from the event horizon.
@@ 2 Dirac equation in Rindler space-time
$ds^{2}=\left(1-v(x)^{2}\right)dt^{2}-2v(x)dtdx-dx^{2},$ (C.6)
where $v(x)$ satisfies $v(x_{h})=\pm 1$, and $v(x)\rightarrow 0$ as $x\rightarrow\infty$ such that the metric of Eq. (C.6) is asymptotically flat. The coordinate transformation $(t,x)\mapsto(\tilde{t},x)$ between Gullstrand-Painlevé coordinates $x^{\mu}=(t,x)$ and Schwarzschild coordinates $\tilde{x}^{\mu}=(\tilde{t},x)$, where
$t(\tilde{t},x)=\tilde{t}+\int dx\frac{v(x)}{1-v(x)^{2}},$ (C.7)
can be used to map the Gullstrand-Painlevé metric line element of Eq. (C.6) to
$ds^{2}=\left(1-v(x)^{2}\right){d\tilde{t}}^{2}-\frac{dx^{2}}{1-v(x)^{2}},$ (C.8)
which is equivalent to the (1+1)-dimensional Schwarzschild metric when the coupling $v(x)$ is given by $v(x)=\sqrt{x_{h}/x}$. This metric contains a coordinate singularity at $v(x_{h})=\pm 1$ and, therefore, it is only valid in the region $x>x_{h}$ that describes the black hole’s exterior.
@@
$f(x)\simeq 2\kappa(x-x_{h})+O(x^{2}),$ (C.9)
where the surface gravity $\kappa=-v^{\prime}(x_{h})$ of the metric’s event horizon has been identified. Hence, near the event horizon of the black hole, Eq. (C.8) for the Schwarzschild metric line element can be expressed as
$ds^{2}=2\kappa(x-x_{h}){d\tilde{t}}^{2}-\frac{dx^{2}}{2\kappa(x-x_{h})}.$ (C.10)
@@
$\rho^{2}=\frac{2(x-x_{h})}{\kappa},$ (C.11)
can be employed to map Eq. (C.10) for the line element of the Schwarzschild metric near the event horizon to
$ds^{2}=(\kappa\rho)^{2}d{\tilde{t}}^{2}-d\rho^{2},$ (C.12)
which is the Rindler metric line element [62, 63] that describes an observer that is uniformly accelerating in Minkowski space-time with a constant proper acceleration $\kappa$. The coordinates $x^{\mu}_{R}=(\tilde{t},\rho)$ will be referred to as Rindler coordinates. The Rindler metric possesses a coordinate singularity at $\rho=0$, which corresponds to the coordinate singularity at $x_{h}$ of the Schwarzschild metric. For this reason, we leave the Rindler metric undefined for the region $\rho<0$ [49]. The zweibein and dual-zweibein corresponding to the Rindler metric line element of Eq. (C.12) are given by
$e^{a}_{\mu}=\begin{pmatrix}\kappa\rho&0\\ 0&1\end{pmatrix},\ \ \ e^{\mu}_{a}=\begin{pmatrix}1/{\kappa\rho}&0\\ 0&1\end{pmatrix}$, (C.13)
respectively.
@@
$(i\gamma^{\mu}D_{\mu}-m)\psi=0,$ (C.14)
$\frac{i}{2}\gamma^{\mu}\partial_{\mu}\psi+\frac{i}{2}\frac{1}{|e|}\partial_{\mu}(|e|\gamma^{\mu}\psi)-\frac{i}{2}\{\gamma^{\mu},\omega_{\mu}\}\psi-m\psi=0$ (C.15)
which, in (1+1)-dimensions, simplifies to
$\frac{i}{2}\gamma^{a}e^{\mu}_{a}\partial_{\mu}\psi+\frac{i}{2}\frac{1}{|e|}\gamma^{a}\partial_{\mu}(|e|e^{\mu}_{a}\psi)-m\psi=0,$ (C.16)
due to the fact that $\{\gamma^{\mu},\omega_{\mu}\}\propto\{\gamma^{a},[\gamma^{b},\gamma^{c}]\}=0$. Expanding out the Einstein summations over $a$ and $\mu$, and substituting in the components of the dual-zweibein of the Rindler metric given in Eq. (C.13), the above Dirac equation can be expressed as
$\begin{split}&\frac{i}{\kappa\rho}\gamma^{0}\partial_{\tilde{t}}\psi(\tilde{t},\rho)+\frac{i}{2}\gamma^{1}\partial_{\rho}\psi(\tilde{t},\rho)\\ &\ \ \ \ \ +\frac{i}{2}\frac{1}{|e|}\gamma^{1}\partial_{\rho}(|e|\psi(\tilde{t},\rho))-m\psi(\tilde{t},\rho)=0\end{split}$ (C.17)
where $|e|=\kappa\rho$. As the Rindler metric is stationary with respect to the Rindler time coordinate $\tilde{t}$, it admits a Killing vector $\zeta^{\mu}=\delta^{\mu}_{\tilde{t}}$ (or equivalently $\partial_{\tilde{t}}$) that is time-like in the region $\rho>0$. Hence, the energy $\epsilon=p_{\tilde{t}}$ of the spinor is a well-defined quantity that is conserved along the Rindler metric’s geodesics [57, 23]. Therefore, the Dirac equation of Eq. (C.17) can be solved via the separation of variables [23, 53, 64]
$\psi(\tilde{t},\rho)=e^{-i\epsilon_{k}\tilde{t}}\psi_{\epsilon_{k}}(\rho).$ (C.18)
Substitution of this solution into the Dirac equation of Eq. (C.17), and making use of the fact that $e^{-i\epsilon_{k}\tilde{t}}\neq 0$ and the identity $(\gamma^{0})^{-1}=-\gamma^{0}$, yields the following eigenvalue equation
$h\psi_{\epsilon_{k}}(\rho)=\epsilon_{k}\psi_{\epsilon_{k}}(\rho),$ (C.19)
@@
$h=i\kappa\rho\gamma^{0}\gamma^{1}\partial_{\rho}+\frac{i}{2}\kappa\gamma^{0}\gamma^{1}-m\kappa\rho\gamma^{0}$ (C.20)
@@
$\langle\phi(\rho),\psi(\rho)\rangle=\int d\mu(\rho)\phi(\rho)^{\dagger}\psi(\rho),$ (C.21)
and $d\mu(\rho)=d\rho$ is a measure of integration. That is, $h$ is Hermitian as it satisfies $\langle\phi,h\psi\rangle=\langle h\phi,\psi\rangle$ [65]. Restricting ourselves to the case of a massless spinor and choosing the chiral representation of the $\gamma$ matrices, $\gamma^{0}=i\sigma^{x}$ and $\gamma^{1}=\sigma^{y}$, the Hermitian operator $h$ can be expressed as
$h=-i\kappa\rho\sigma^{z}\partial_{\rho}-\frac{i}{2}\kappa\sigma^{z}.$ (C.22)
This allows the eigenvalue equation given in Eq. (C.20) to be expressed as the two following decoupled differential equations:
$\epsilon_{k}\psi^{\pm}_{\epsilon_{k}}(\rho)\pm i\kappa\rho\partial_{\rho}\psi^{\pm}_{\epsilon_{k}}(\rho)\pm\frac{i\kappa}{2}\psi^{\pm}_{\epsilon_{k}}(\rho)=0,$ (C.23)
@@
$\psi_{\epsilon_{k}}(\rho)=\frac{1}{\sqrt{2\pi|\kappa|}}\begin{pmatrix}\rho^{\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}\\ \rho^{-\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}\end{pmatrix}$. (C.24)
@@
$\phi_{\epsilon_{k}}(\rho)=\psi_{\epsilon_{k}}^{*}(\rho)=\frac{1}{\sqrt{2\pi|\kappa|}}\begin{pmatrix}\rho^{-\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}\\ \rho^{\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}\end{pmatrix}$. (C.25)
Analogous to $\psi_{\epsilon_{k}}(\rho)$, $\phi_{\epsilon_{k}}(\rho)$ can be expressed via the linear combination $\phi_{\epsilon_{k}}(\rho)=\varphi^{+}_{\epsilon_{k}}(\rho)+\varphi^{-}_{\epsilon_{k}}(\rho)$ of two orthonormal spinors $\varphi^{\pm}_{\epsilon_{k}}(\rho)=\mathcal{N}\rho^{\mp\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}u_{\pm}$ that solve the Dirac equation and satisfy the eigenvalue equation $h\varphi^{\pm}_{\epsilon_{k}}(\rho)=-\epsilon_{k}\varphi^{\pm}_{\epsilon_{k}}(\rho)$. A general solution to the Dirac equation must be constructed with $\psi_{\epsilon_{k}}$ and $\phi_{\epsilon_{k}}$ via an eigenfunction decomposition.
Assuming that the orthonormal basis formed by the eigenfunctions $\psi_{\epsilon_{k}}(\rho)$ of the Hermitian operator $h$ is complete, any function $F(\rho)$ within the Hilbert space $\mathcal{H}_{R}$ defined by the inner product of Eq. (C.21) can be decomposed into a continuous sum of the eigenfunctions $\psi_{\epsilon_{k}}(\rho)$ as [53, 64, 65]
$F(\rho)=\int d\mu(k)\tilde{f}(k)\psi_{\epsilon_{k}}(\rho),$ (C.26)
@@
$\langle\tilde{f}(k),\tilde{g}(k)\rangle=\int d\mu(k)\tilde{f}^{\dagger}(k)\tilde{g}(k)$ (C.27)
of a corresponding Hilbert space $\mathcal{L}^{2}(k)$ of functions $\tilde{f}(k)$, which is related to $\mathcal{H}_{R}$ via the eigenfunction decomposition of Eq. (C.26) and its inverse [53, 64, 65]
$\tilde{f}(k)=\int d\mu(\rho)\psi^{\dagger}_{\epsilon_{k}}(\rho)F(\rho).$ (C.28)
@@
Using the eigenfunction decomposition defined in Eq. (C.26), the most general solution $\psi(\tilde{t},\rho)$ of the Dirac equation in a Rindler space-time can be decomposed into the continuous sum of both the spinor $\psi_{\epsilon_{k}}(\rho)$ and charge conjugate spinor $\phi_{\epsilon_{k}}(\rho)$ eigenfunctions of the Dirac equation. That is,
$\psi(\tilde{t},\rho)=\int^{\infty}_{0}d\mu(k)\left(\tilde{f}(k)e^{-i\epsilon_{k}\tilde{t}}\psi_{\epsilon_{k}}(\rho)+\tilde{g}(k)e^{i\epsilon_{k}\tilde{t}}\phi_{\epsilon_{k}}(\rho)\right).$ (C.29)
From the eigenvalue equation $h\psi_{\epsilon_{k}}(\rho)=\epsilon_{k}\psi_{\epsilon_{k}}(\rho)$, it is evident that $\psi_{\epsilon_{k}}(\rho)$ is a positive energy solution when $\epsilon_{k}>0$ and a negative energy solution when $\epsilon_{k}<0$. Similarly, from $h\phi_{\epsilon_{k}}(\rho)=-\epsilon_{k}\phi_{\epsilon_{k}}(\rho)$, it is evident that $\phi_{\epsilon_{k}}(\rho)$ is a positive energy solution when $\epsilon_{k}<0$ and a negative energy solution when $\epsilon_{k}>0$. To ensure that the spinor $\psi_{\epsilon_{k}}(\rho)$ and its charge conjugate $\phi_{\epsilon_{k}}(\rho)$ always correspond to positive and negative energy solutions, respectively, we define
$k=\begin{cases}\epsilon_{k},&\text{if }\epsilon_{k}\geq 0\\ -\epsilon_{k},&\text{if }\epsilon_{k}<0\end{cases},\ \text{and}\ u_{k}=\begin{cases}u_{+},&\text{if }k\geq 0\\ u_{-},&\text{if }k<0\end{cases},$ (C.30)
such that the general solution to the Dirac equation in Eq. (C.29) can be expressed as [51, 49]
$\psi(\tilde{t},\rho)=\int^{\infty}_{-\infty}d\mu(k)\left(\tilde{f}(k)e^{-i|k|\tilde{t}}\psi_{k}(\rho)+\tilde{g}(k)e^{i|k|\tilde{t}}\phi_{k}(\rho)\right),$ (C.31)
@@
$\begin{split}h\psi_{k}(\rho)=\begin{cases}k\psi_{k}(\rho),&\text{for }k\geq 0\\ -k\psi_{k}(\rho),&\text{for }k<0\end{cases},\\ h\phi_{k}(\rho)=\begin{cases}-k\phi_{k}(\rho),&\text{for }k\geq 0\\ k\phi_{k}(\rho),&\text{for }k<0\end{cases}.\end{split}$ (C.32)

### v2

@@ 1 Surface gravity of Gullstrand-Painlevé metric
Consider the (1+1)-dimensional Gullstrand-Painlevé metric line element given in Eq. (10), which is the line element of the (1+1)-dimensional Schwarzschild metric expressed in Gullstrand-Painlevé coordinates $x^{\mu}=(t,x)$. Here, the couplings $u$ and $v$ are independent of $t$, but not necessarily of $x$; thus, the metric is stationary, though not static [66]. As the Gullstrand-Painlevé metric is stationary, it admits a Killing vector $\xi^{\mu}=\delta^{\mu}_{t}=(1,0)$ [48], which is time-like in the region $u^{2}>v^{2}$ and satisfies the Killing equation
$\xi_{\mu;\nu}+\xi_{\nu;\mu}=0.$ (B.1)
The vector $\xi_{\mu}=g_{\mu\nu}\xi^{\nu}=(1-v^{2}/u^{2},-v/u^{2})$ is also a Killing vector that satisfies the Killing equation. As the Gullstrand-Painlevé metric is regular at the event horizon $x_{h}$, these Killing vectors can be used to find the surface gravity $\kappa$ of the black hole space-time, as defined by a distant observer, by evaluating
$\xi^{\mu}\xi_{\nu;\mu}=\xi^{\mu}\left(\xi_{\nu,\mu}-\xi_{\gamma}\tilde{\Gamma}^{\gamma}_{\nu\mu}\right)=\kappa\xi_{\nu},$ (B.2)
at the black hole’s event horizon [69, 68].
Expanding the Einstein summations over $\mu$ and $\gamma$, and making use of the fact that $\xi_{\nu,t}=\forall\nu$, $\xi^{t}=1$, and $\xi^{x}=0$, Eq. (B.2) for the surface gravity can be expressed as
$\xi_{x}\tilde{\Gamma}^{x}_{\nu t}-\xi_{t}\tilde{\Gamma}^{t}_{\nu t}=\kappa\xi_{\nu},$ (B.3)
@@
$\frac{1}{2}\xi_{t}g^{tt}g_{tt,x}-\frac{1}{2}\xi_{x}g^{xt}g_{tt,x}=\kappa\xi_{x}.$ (B.4)
Substitution of the relevant values for the components of the metric tensor corresponding to the Gullstrand-Painlevé line element of Eq. (10) into the above equation and evaluating the resultant expression at the event horizon gives the following surface gravity:
$\kappa=\pm\frac{du(x_{h})}{dx}-\frac{dv(x_{h})}{dx}.$ (B.5)
@@ 2 Dirac equation in Rindler space-time
$ds^{2}=\left(1-v(x)^{2}\right)dt^{2}-2v(x)dtdx-dx^{2},$ (B.6)
where $v(x)$ satisfies $v(x_{h})=\pm 1$, and $v(x)\rightarrow 0$ as $x\rightarrow\infty$ such that the metric of Eq. (B.6) is asymptotically flat. The coordinate transformation $(t,x)\mapsto(\tilde{t},x)$ between Gullstrand-Painlevé coordinates $x^{\mu}=(t,x)$ and Schwarzschild coordinates $\tilde{x}^{\mu}=(\tilde{t},x)$, where
$t(\tilde{t},x)=\tilde{t}+\int dx\frac{v(x)}{1-v(x)^{2}},$ (B.7)
can be used to map the Gullstrand-Painlevé metric line element of Eq. (B.6) to
$ds^{2}=\left(1-v(x)^{2}\right){d\tilde{t}}^{2}-\frac{dx^{2}}{1-v(x)^{2}},$ (B.8)
which is equivalent to the (1+1)-dimensional Schwarzschild metric when $v(x)=\sqrt{x_{h}/x}$. The metric of Eq (B.8) contains a coordinate singularity at $v(x_{h})=\pm 1$ and is therefore only valid in the region $x>x_{h}$ that describes the black hole’s exterior.
@@
$f(x)\simeq 2\kappa(x-x_{h})+O(x^{2}),$ (B.9)
where the metric’s surface gravity $\kappa=-v^{\prime}(x_{h})$ has been identified. Hence, near the event horizon of the black hole, Eq. (B.8) for the Schwarzschild metric line element can be expressed as
$ds^{2}=2\kappa(x-x_{h}){d\tilde{t}}^{2}-\frac{dx^{2}}{2\kappa(x-x_{h})}.$ (B.10)
@@
$\rho^{2}=\frac{2(x-x_{h})}{\kappa},$ (B.11)
can be employed to map Eq. (B.10) for the line element of the Schwarzschild metric near the horizon to
$ds^{2}=(\kappa\rho)^{2}d{\tilde{t}}^{2}-d\rho^{2},$ (B.12)
which is the Rindler metric line element [53, 54] that describes an observer that is uniformly accelerating in Minkowski space-time with a constant proper acceleration $\kappa$. The coordinates $x^{\mu}_{R}=(\tilde{t},\rho)$ will be referred to as Rindler coordinates. The Rindler metric possesses a coordinate singularity at $\rho=0$, which corresponds to the coordinate singularity at $x_{h}$ of the Schwarzschild metric. For this reason, we leave the Rindler metric undefined for the region $\rho<0$ [41]. The zweibein corresponding to the Rindler metric line element of Eq. (B.12) is given by $e^{a}_{\mu}=\text{diag}(\kappa\rho,\ 1)$.
@@
$(i\gamma^{\mu}D_{\mu}-m)\psi=0,$ (B.13)
$\frac{i}{2}\gamma^{\mu}\partial_{\mu}\psi+\frac{i}{2}\frac{1}{|e|}\partial_{\mu}(|e|\gamma^{\mu}\psi)-\frac{i}{2}\{\gamma^{\mu},\omega_{\mu}\}\psi-m\psi=0,$ (B.14)
which in (1+1)-dimensions simplifies to
$\frac{i}{2}\gamma^{a}e^{\mu}_{a}\partial_{\mu}\psi+\frac{i}{2}\frac{1}{|e|}\gamma^{a}\partial_{\mu}(|e|e^{\mu}_{a}\psi)-m\psi=0$ (B.15)
due to the fact that $\{\gamma^{\mu},\omega_{\mu}\}\propto\{\gamma^{a},[\gamma^{b},\gamma^{c}]\}=0$. As the Rindler metric is stationary with respect to the Rindler time coordinate $\tilde{t}$, it admits a Killing vector $\zeta^{\mu}=\delta^{\mu}_{\tilde{t}}$ (or equivalently $\partial_{\tilde{t}}$) that is time-like in the region $\rho>0$. Hence, the energy $\epsilon=p_{\tilde{t}}$ of the spinor is a well-defined quantity that is conserved along the Rindler metric’s geodesics [48, 35]. Therefore, the Dirac equation can be solved via the separation of variables [35, 23, 25]
$\psi(\tilde{t},\rho)=e^{-i\epsilon_{k}\tilde{t}}\psi_{\epsilon_{k}}(\rho).$ (B.16)
Substitution of this solution, along with the components of the Rindler metric zweibein, into the Dirac equation of Eq. (B.15) yields the following eigenvalue equation
$h_{R}\psi_{\epsilon_{k}}(\rho)=\epsilon_{k}\psi_{\epsilon_{k}}(\rho),$ (B.17)
@@
$h_{R}=i\kappa\rho\gamma^{0}\gamma^{1}\partial_{\rho}+\frac{i}{2}\kappa\gamma^{0}\gamma^{1}-m\kappa\rho\gamma^{0}$ (B.18)
@@
$\langle\phi(\rho),\psi(\rho)\rangle=\int d\mu(\rho)\phi(\rho)^{\dagger}\psi(\rho),$ (B.19)
and $d\mu(\rho)=d\rho$ is a measure of integration. That is, $h_{R}$ is Hermitian as it satisfies $\langle\phi,h_{R}\psi\rangle=\langle h_{R}\phi,\psi\rangle$ [24]. Restricting ourselves to the case of a massless spinor and choosing the chiral representation of the $\gamma$ matrices, $\gamma^{0}=i\sigma^{x}$ and $\gamma^{1}=\sigma^{y}$, the Hermitian operator $h_{R}$ can be expressed as
$h_{R}=-i\kappa\rho\sigma^{z}\partial_{\rho}-\frac{i}{2}\kappa\sigma^{z}.$ (B.20)
This allows the eigenvalue equation given in Eq. (B.17) to be expressed as the two following decoupled differential equations:
$\epsilon_{k}\psi^{\pm}_{\epsilon_{k}}(\rho)\pm i\kappa\rho\partial_{\rho}\psi^{\pm}_{\epsilon_{k}}(\rho)\pm\frac{i\kappa}{2}\psi^{\pm}_{\epsilon_{k}}(\rho)=0,$ (B.21)
@@
$\psi_{\epsilon_{k}}(\rho)=\frac{1}{\sqrt{2\pi|\kappa|}}\begin{pmatrix}\rho^{\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}\\ \rho^{-\frac{i\epsilon_{k}}{\kappa}-\frac{1}{2}}\end{pmatrix}$. (B.22)
@@
Assuming that the orthonormal basis formed by the eigenfunctions $\psi_{\epsilon_{k}}(\rho)$ of the Hermitian operator $h_{R}$ is complete, any function $F(\rho)$ within the Hilbert space $\mathcal{H}_{R}$ defined by the inner product of Eq. (B.19) can be decomposed into a continuous sum of the eigenfunctions $\psi_{\epsilon_{k}}(\rho)$ as [23, 25, 24]
$F(\rho)=\int d\mu(k)\tilde{f}(k)\psi_{\epsilon_{k}}(\rho),$ (B.23)
@@
$\langle\tilde{f}(k),\tilde{g}(k)\rangle=\int d\mu(k)\tilde{f}^{\dagger}(k)\tilde{g}(k)$ (B.24)
of a corresponding Hilbert space $\mathcal{L}^{2}(k)$ of functions $\tilde{f}(k)$, which is related to $\mathcal{H}_{R}$ via the eigenfunction decomposition of Eq. (B.23) and its inverse [23, 25, 24]
$\tilde{f}(k)=\int d\mu(\rho)\psi^{\dagger}_{\epsilon_{k}}(\rho)F(\rho).$ (B.25)
@@
Using the eigenfunction decomposition defined in Eq. (B.23), the most general solution $\psi(\tilde{t},\rho)$ to the Dirac equation in Rindler space-time can be decomposed into a continuous sum of the Dirac equation’s spinor $\psi_{\epsilon_{k}}(\rho)$ and charge conjugate spinor $\phi_{\epsilon_{k}}(\rho)$ eigenfunctions. That is,
$\psi(\tilde{t},\rho)=\int^{\infty}_{0}d\mu(k)\left(\tilde{f}(k)e^{-i\epsilon_{k}\tilde{t}}\psi_{\epsilon_{k}}(\rho)+\tilde{g}(k)e^{i\epsilon_{k}\tilde{t}}\phi_{\epsilon_{k}}(\rho)\right).$ (B.26)
From the eigenvalue equation of Eq. (B.17), it is evident that $\psi_{\epsilon_{k}}(\rho)$ is a positive energy solution when $\epsilon_{k}>0$ and a negative energy solution when $\epsilon_{k}<0$, with the converse being true for $\phi_{\epsilon_{k}}(\rho)$. Thus, to ensure that the spinor $\psi_{\epsilon_{k}}(\rho)$ and its charge conjugate $\phi_{\epsilon_{k}}(\rho)$ always correspond to positive and negative energy solutions, respectively, we define
$k=\begin{cases}\epsilon_{k},&\text{if }\epsilon_{k}\geq 0\\ -\epsilon_{k},&\text{if }\epsilon_{k}<0\end{cases},\ \text{and}\ u_{k}=\begin{cases}u_{+},&\text{if }k\geq 0\\ u_{-},&\text{if }k<0\end{cases},$ (B.27)
such that the general solution to the Dirac equation in Eq. (B.26) can be expressed as [20, 41]
$\psi(\tilde{t},\rho)=\int^{\infty}_{-\infty}d\mu(k)\left(\tilde{f}(k)e^{-i|k|\tilde{t}}\psi_{k}(\rho)+\tilde{g}(k)e^{i|k|\tilde{t}}\phi_{k}(\rho)\right),$ (B.28)
@@
$\begin{split}h_{R}\psi_{k}(\rho)=\begin{cases}k\psi_{k}(\rho),&\text{for }k\geq 0\\ -k\psi_{k}(\rho),&\text{for }k<0\end{cases},\\ h_{R}\phi_{k}(\rho)=\begin{cases}-k\phi_{k}(\rho),&\text{for }k\geq 0\\ k\phi_{k}(\rho),&\text{for }k<0\end{cases}.\end{split}$ (B.29)